OpenStudy (anonymous):

A mechanic repairs Mr. Estes car. The amount for parts is $48 and the rate for the mechanic is $40 per hour. Write and solve an equation to find the total cost of repairs to Mr. Estes car if the mechanic works for 1.5 hours.

OpenStudy (anonymous):

40+40+20=100 100+48=148 answer-$148

OpenStudy (anonymous):

How did you get two 40's?

OpenStudy (anonymous):

oh gosh sorry! sorry sorry okay 40+20= 60 60+48=108 $108

OpenStudy (anonymous):

That's what I got the first time. Thank you (:

OpenStudy (anonymous):

your welcome sorry for the mistake . i feel bad

OpenStudy (anonymous):

:)